V = 8Vcos (4+- 2007 {Rule I GME find CR & CAR AND cci Val
Answer #1

Solution: R=32 ZR2 TOME I WL circuit in Time domain & lt) = scosllyt 20 V source voltage in phasor foor vs = 813 w = 4 radlec(0.5823 +9 2.4805) Na = 8128 8.5833) 8.008 va =81-28 Va = 81-25 va = 13.7151-20.00) = 137151-28 v. correst Brough he is IR =